OpenStudy (anonymous):

Romero has the following scores on his math tests: 79, 86, 95, 88. How much would his mean change if the lowest test score is dropped? Round to the nearest whole number. A) 6 B)5 C) 4 D) 3

OpenStudy (anonymous):

Mean equals average. To get the average you sum all the values and divide by the number of values you have. So: 79+86+95+88= 348 348/4 (since you have 4 values) = 87 By dropping the lowest score, 79, you would have: 86+95+88 = 269 269/3 (because now you have 3 values) = 89.666 therefore the change is a difference between 89.666 and 87: 89.666- 87 = 2.666, roughly 3. Hows that?

OpenStudy (anonymous):

Really Helped!!! Thanks!

OpenStudy (anonymous):

You're welcome